SYMMETRY PECULIARITIES OF THE INTRACRYSTALLINE FIELDS LAYERED SEMICONDUCTOR CRYSTALS (PbI_{2})_{(1−x)} (BiI_{3})_{(x)}

\underline{IGOR VERTEGEL}, EUGENE CHESNOKOV and ALEKSANDR OVCHARENKO, Institute of Physics National Academy of Sciences of Ukraine 46, Prospect Nauki, 03680 Kiev, Ukraine.

In this work the results of the investigation of the ^{127}I NQR spectra at 77K for mixed layered semiconductor crystals (PbI_{2})_{(1−x)} (BiI_{3})_{(x)} in a wide range of value (0 <x <0.50) are presented. It is shown that in the range 0.05 <x <0.20 of admixture PbI_{2} the observed behavior of parameters of the ^{127}I NQR spectra testify about entrance of admixture atoms PbI_{2} into the crystal layers. It is shown, that at 0.05 < x <0.20 clusters from groups of atoms PbI_{2} insular type can be formed, which lay within the limits of the layers of crystal (PbI_{2})_{(1−x)} (BiI_{3})_{(x)}. Upon further increasing of the containing of admixture PbI_{2} in crystal BiI_{3} the new ^{127}I NQR line is appearing. The observed at x_{}˜ 0.20 the new line in spectrum 127I NQR can testify that the mixed crystal (PbI_{2})_{(1−x)} (BiI_{3})_{(x)} undergoes structural phase transition. It is concluded that at x>0.20 a new crystal presents a solid mixture glassy crystal of substitution type in which of PbI_{2} atoms are fully or partially are ordering and lay between crystal^{a} .

aA.I.Barabash, I.G.Vertegel, E.D.Chesnokov et.al., Ukr. J.Phys., 2011,vol.56, No.2, p.158-160.